# Program Synthesis Dialog Agents for Interactive Decision-Making

### v2

@@ 5 Experimental Results
Table 1: F1 score and dialog turns to completion for ProADA and baseline models.
@@ 6 Failure Analysis
We observe multiple distinct types of errors that contribute to poor reasoning and inefficient dialog. Program synthesis-guided dialog reduces errors overall, but introduces unique failure modes associated with code generation.
6.1 General Errors
Several failure modes persist across all strategies, indicating core weaknesses in foundational model reasoning ability.
@@ 6.1 General Errors
Domain knowledge & edge cases: Models fail to account for uncommon edge cases, such as children living with parents but not being claimed as tax dependents.
6.2 Baseline Errors
Although it is difficult to confidently attribute final predictions to specific mistakes during question generation, we observe several flawed reasoning patterns when using direct and ReAct prompting:
Hallucination: Baseline models frequently return True in Ready before collecting all relevant information, implying either a logical reasoning failure or a hallucination of relevant facts.
Ultra-specificity: Models ask needlessly specific questions ("Is your total annual income below \$69,900?") when a more general question ("What is your total annual investment income?") would produce information useful elsewhere, resulting in superfluous dialog turns.
@@ 6.2 Baseline Errors
Conflating users: Baseline models sometimes conflate household members or fail to specify which member they are asking about.
6.3 ProADA Errors
@@
No recovery: Since ProADA models strictly adhere to dialog planning according to the generated code, they have no ability to recover or deviate when users generate an unexpected response, such as "I don’t know."
@@ 6.4 Errors by Simulated Users
Authors annotated 61 simulated user responses for faithfulness to the user profile, finding 60 (98.4%) of questions are answered faithfully. The simulated user tends towards verbosity, providing additional unrequested information in 5 cases (8.2%). We find unnatural but faithful responses in 2 cases (3. 3%), indicating that the frequency of errors due to simulated user misbehavior is low. In qualitative probing, we find that the simulated user can respond accurately to diverse questions up to two hops (e.g., "How many children do you have under the age of 5?"). Sufficiently complex queries or those with more than two hops tend to cause the simulated user to respond that it cannot answer the question, but we rarely observe models generating such questions in our experiments.
@@ 7 Discussion
Synthesized code also serves as a window into the agent’s decisions. Although we generate code automatically in this work, the code may be checked manually or with software tools to ensure correctness before deployment. Unlike black-box models, program synthesis-guided models like ProADA may also be subject to unit tests to ensure code quality.
@@ 8 Related Work
Many dialog agent tasks have been proposed, including offline task-oriented dialog Andreas et al. (2020) Budzianowski et al. (2018) and online user simulations using real humans or LM agents as responders Gür et al. (2018) He et al. (2018). Question generation is a related task where agents seek information relevant to a downstream task, such as user intent Min et al. (2020) or relevant facts Toles et al. (2023). Some task-oriented dialog datasets focus on clarification and information seeking, such as Zhang et al. (2023). However, datasets such as ShARC Saeidi et al. (2018) and ClariT Feng et al. (2023) only require "yes" or "no" questions. BeNYfits expands on these works by adding a highly realistic, multi-turn dialog agent task requiring logical reasoning and domain-specific knowledge. Similar tasks include MediQ Li et al. (2024), which benchmarks medical diagnosis through dialog, and ClarQ-LLM Gan et al. (2024), which focuses on discovering hidden information while playing an adventurer. In comparison, BeNYfits focuses on logically reasoning legalistic tasks to reach a binary prediction.
@@ 9 Conclusion
We present a strong tool-augmented method to solve interactive decision-making in dialogs and a novel and realistic benchmark for measuring decision-problem accuracy and dialog efficiency. Our method ameliorates memory and planning issues by converting key information in user utterances into structured key-value pairs to improve reasoning, latency, and cost by offloading computations onto an agent-created Python tool. Such structured coding support overcomes many problems of pure LLM baselines such as hallucination of missing information, lack of object tracking, being over-confident, etc. Ultimately, our proposed method achieved an F1 score of 55.6 (compared to at most 42.2 for the baselines) while reducing the dialog turns needed by more than 30% compared to the next best agent, raising hopes for reducing user burden and increasing access to public opportunities using language models.

### v3

@@ 5.1 User Study
Table 1: F1 score and dialog turns to completion for ProADA and baseline models. Average is averaged across Diverse and Representative Datasets. Weighted F1 score uses equation (1).
@@ 6 Failure Analysis
We observe multiple distinct types of errors that contribute to poor reasoning and inefficient dialog. Program synthesis-guided dialog reduces errors overall, but introduces unique failure modes associated with code generation. However, several failure modes persist across all strategies, indicating core weaknesses in foundational model reasoning ability.
@@
Domain knowledge & edge cases: Models fail to account for edge cases, such as 17-year-olds with work income or adult dependents.
6.1 Baseline Behavior
Although it is difficult to confidently attribute final predictions to specific mistakes in black-box models during question generation, we observe several flawed reasoning patterns when using direct and ReAct prompting:
Hallucination: Baseline models frequently return True in Ready before collecting all relevant information, implying either a logical reasoning failure or an internal hallucination of relevant facts.
Hyperspecificity: Models ask needlessly specific questions ("Is your total annual income below \$69,900?") when a more general question ("What is your total annual investment income?") would produce information useful elsewhere, resulting in superfluous dialog turns.
@@ 6.1 Baseline Behavior
Conflating Users: Baseline models often conflate household members or fail to specify which member they are asking about.
6.2 ProADA Behavior
@@
To investigate bias arising from the mapping of user responses to structured data, we conduct the following study: We analyze 17 questions generated by ProADA and provide 6 perturbed versions of the same answer. Across all 102 trials, we find that overall, user responses are rarely mapped incorrectly, in only 2% (n=2) of cases.
[!htb] ProADA GPT-4o + React Accuracy ↑ 81.3% 42.1% Discovered Opportunity ↑ 80.0% 50.0% Clarity ↑ 4.80 4.14 Ease of Use ↑ 4.53* 3.96 Could Save Time ↑ 4.47 3.39 Question Appropriateness ↑ 4.43 3.54 Task Complexity ↓ 1.73 1.89 Workload ↓ 1.33 2.04 Irritation ↓ 1.50 2.71 User-reported accuracy, usability, and frequency of discovery of at least one opportunity for which the user is eligible. * p<0.05, † p<0.01
@@ 6.3 Errors by Simulated Users
Authors annotated 61 simulated user responses for faithfulness to the user profile, finding 60 (98.4%) of questions are answered faithfully (Figure 5). The simulated user tends towards verbosity, providing additional unrequested information in 5 cases (8.2%). We find unnatural but faithful responses in 2 cases (3. 3%), indicating that the frequency of errors due to simulated user misbehavior is low. In qualitative probing, we find that the simulated user can respond accurately to diverse questions up to two hops (e.g., "How many children do you have under the age of 5?"). Sufficiently complex queries or those with more than two hops tend to cause the simulated user to respond that it cannot answer the question, but we rarely observe models generating such questions in our experiments.
@@ 7 Discussion
This white-box reasoning approach enables transparency into model failure modes and clear opportunities for future work. On the dialog side, we observe that LLMs often get stuck in dialog loops after making false assumptions about the user. In code generation, we observe a significant portion of errors are due to inaccurate code generation (Figure 5) related to domain expertise. Specifically, coding agents could search for domain knowledge (e.g., the tax-purposes definition of a dependent in the state of New York) before conducting program synthesis. Strategies in self-correction and theory of mind may be applicable here. Although we generate code automatically in this work, the code may be checked manually or with software tools to ensure correctness before deployment. Unlike black-box models, program synthesis-guided models like ProADA may also be subject to unit tests to ensure code quality.
@@ 8 Related Work
Many dialog agent tasks have been proposed, including offline task-oriented dialog Andreas et al. (2020) Budzianowski et al. (2018) and online user simulations using real humans or LM agents as responders Gür et al. (2018) He et al. (2018). Question generation is a related task where agents seek information relevant to a downstream task, such as user intent Min et al. (2020), relevant facts Toles et al. (2023), or user preferences Li et al. (2023). Some task-oriented dialog datasets focus on clarification and information seeking, such as Zhang et al. (2023). However, datasets such as ShARC Saeidi et al. (2018) and ClariT Feng et al. (2023) only require "yes" or "no" questions. BeNYfits expands on these works by adding a highly realistic, multi-turn dialog agent task requiring logical reasoning and domain-specific knowledge. Similar tasks include MediQ Li et al. (2024), which benchmarks medical diagnosis through dialog, and ClarQ-LLM Gan et al. (2024), which focuses on discovering hidden information while playing an adventurer. In comparison, BeNYfits focuses on logically reasoning legalistic tasks to reach a binary prediction.
@@ 9 Conclusion
We present a strong tool-augmented method to solve interactive decision-making in dialogs and a novel and realistic benchmark for measuring decision-problem accuracy and dialog efficiency. Our method ameliorates memory and planning issues by converting key information in user utterances into structured key-value pairs to improve reasoning, latency, and cost by offloading computations onto an agent-created Python tool. Such structured coding support overcomes many problems of pure LLM baselines such as hallucination of missing information, lack of object tracking, being over-confident, etc. Ultimately, our proposed method achieved an F1 score of 56.2 (compared to at most 42.2 for the baselines) while reducing the dialog turns needed by 13.9% compared to the next best agent, raising hopes for reducing user burden and increasing access to public opportunities using language models.
@@ 10 Limitations
Although our dataset includes numerous state and federal level opportunities, suggesting broader applicability to many US-oriented applications, we agree that NYC Open Data may not generalize to other regions and contexts. We will add the following to the limitations section:
We note that, due to its size and budget, New York City and State provide a wider range of services than most other municipalities around the world. For example, New York State’s right to shelter drives many of the 18 housing-related opportunities in the dataset. Other domains, especially those requiring subjective evaluation (e.g., merit-based opportunities) present additional challenges, as eligibility cannot necessarily be determined consistently based on self-reported user information.
@@ 11 Ethical Considerations
Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. (DGE-2036197). We thank Nick Deas and Zachary Horvitz for their helpful feedback on early drafts.
Appendix A Structured Data Mapping
To investigate bias arising from the mapping of user responses to structured data, we conduct the following study: We analyze 17 questions generated by ProADA and provide 6 perturbed versions of the same answer. For example, to the question “What is the number of people living in your household?”, we provide the following perturbations:
Numeric: 1
Text: one
Verbose: There is only one person in my household
Multi-hop: I don’t live with anyone else
Misspelled: onee
Extraneous Info: One but I have a dog
Across all 102 trials, we find that overall, user responses are mapped incorrectly only 2% (n=2) of the time. Inaccurate mapping occurred only once each in multi-hop and extraneous info perturbations. Otherwise, we see ProADA attempt to clarify user responses in 6% (n=1) of verbose perturbations and 23% (n=4) misspelling perturbations, which we consider desired behavior. We believe two factors contribute to the low level of bias observed: Firstly, eligibility requirements in BeNYfits are objective, leaving little room for biased interpretation. Secondly, ProADA generates primarily yes/no (65%, n=11) and numeric questions (35%, n=6), which are generally straightforward to parse.
Appendix B Detailed Results
Diversity Representative Average
Strategy Model Accuracy Recall Precision F1 Turns Accuracy Recall Precision F1 Turns F1 Turns TW F1
Direct Prompting Llama 3.1 8B 51.5 37.7 49.1 42.6 8.92 63.8 41.7 19.0 26.1 5.01 34.4 7.0 32.1
51.5 37.7 49.1 42.6 3.85 65.1 22.2 12.9 16.3 10.00 29.5 6.9 27.6
55.7 43.8 54.7 48.7 3.69 66.8 36.1 19.1 25.0 4.32 36.8 4.0 35.4
52.9 39.7 51.0 44.6 5.49 65.2 33.3 17.0 22.5 6.44 33.6 6.0 31.7
Llama 3.1 70B 52.1 53.4 50.0 51.7 46.75 66.8 52.8 23.8 32.8 76.08 42.2 61.4 26.1
53.8 63.0 51.4 56.6 37.05 66.0 41.7 20.3 27.3 84.08 41.9 60.6 26.1
53.4 58.9 51.2 54.8 41.07 62.6 33.3 15.8 21.4 85.08 38.1 63.1 23.4
53.1 58.4 50.9 54.3 41.62 65.1 42.6 19.9 27.2 81.75 40.8 61.7 25.2
GPT-4o 77.4 33.3 29.3 31.2 16.75 55.1 38.4 54.4 45.0 37.64 38.1 27.2 29.9
GPT-o3 mini 82.6 63.9 45.1 52.9 20.32 77.0 13.9 17.9 15.6 31.20 34.2 25.8 27.2
Claude 3.7 54.4 37.0 53.5 43.7 11.91 76.6 16.7 19.4 17.9 8.88 30.8 10.4 27.9
ReAct Agent Llama 3.1 8B 56.2 41.7 15.5 22.6 13.96 56.2 41.7 15.5 22.6 13.96 22.6 14.0 19.8
50.8 49.3 48.6 49.0 6.07 56.6 47.2 17.0 25.0 8.52 37.0 7.3 34.5
52.1 49.3 50.0 49.7 8.38 63.4 52.8 21.6 30.6 10.36 40.2 9.4 36.7
53.0 46.8 38.0 40.4 9.47 58.7 47.2 18.0 26.1 10.95 33.2 10.2 30.2
Llama 3.1 70B 57.4 62.3 54.8 58.3 9.53 63.8 41.7 19.0 26.1 37.92 42.2 23.7 34.1
GPT-4o 56.7 45.9 55.8 50.4 18.41 71.1 25.0 18.0 20.9 13.28 35.7 15.8 30.8
Claude 3.7 57.6 31.0 61.6 41.3 38.66 77.9 13.9 20.0 16.4 28.44 28.8 33.6 21.6
ProADA w/o Clarity Llama 3.1 8B 62.6 41.8 67.8 51.7 16.58 75.7 58.3 33.3 42.4 22.96 47.1 19.8 39.3
61.6 43.8 64.6 52.2 18.14 78.7 61.1 37.9 46.8 27.96 49.5 23.1 40.2
59.2 39.0 62.0 47.9 15.58 76.2 44.4 30.8 36.4 19.24 42.1 17.4 35.9
61.2 41.6 64.8 50.6 16.77 76.9 54.6 34.0 41.9 23.39 46.2 20.1 38.5
Llama 3.1 70B 64.9 49.3 68.6 57.4 18.83 86.0 72.2 53.1 61.2 19.12 59.3 19.0 49.8
63.0 43.8 67.4 53.1 18.55 83.0 61.1 45.8 52.4 20.20 52.7 19.4 44.2
61.5 42.5 65.3 51.5 17.28 78.3 38.9 32.6 35.4 20.16 43.4 18.7 36.6
63.1 45.2 67.1 54.0 18.22 82.4 57.4 43.8 49.7 19.83 51.8 19.0 43.5
GPT-4o 63.3 39.0 71.3 50.4 15.73 86.8 66.7 55.8 60.8 17.20 55.6 16.5 47.7
ProADA Llama 3.1 70B 70.2 53.4 77.2 63.2 20.32 82.6 63.9 45.1 52.9 24.20 58.0 22.3 47.5
66.6 54.8 69.0 61.1 20.38 84.7 55.6 50.0 52.6 24.80 56.9 22.6 46.4
67.2 52.1 71.7 60.3 18.75 80.0 61.1 40.0 48.4 22.04 54.3 20.4 45.1
66.8 49.8 72.3 58.7 18.79 83.5 61.8 47.7 53.7 22.06 56.2 20.4 46.7
Claude 3.7 (code) 67.9 44.5 79.3 57.0 21.5 87.7 55.6 60.6 58.0 22.44 57.5 22.0 47.1
Qwen 32B (code) 53.1 13.7 54.1 21.9 10.1 76.6 16.7 19.4 17.9 11.52 19.9 10.8 17.9
Table 2: Complete BeNYfits results
Appendix C List of Prompts
C.1 “Are Benefits Ready?” Prompt
Eligibility requirements: {eligibility_requirements}.
Is the information sufficient to determine whether any member of the user’s household is eligible for all programs? Answer only in one word True or False.
C.2 “Predict Benefits Eligibility” Prompt
Eligibility: {eligibility_requirements}.
Predict the programs for which any member of the user’s household is eligible. Return only a boolean array of length {num_programs}, e.g. {example_array}, where the value at index ‘i‘ is true iff the user is eligible for program ‘i‘. Only return the array. Do not return anything else in the response. If a user’s eligibility is unclear, make your best guess.
C.3 “Ask a Clarifying Question” Prompt
Eligibility: {eligibility_requirements}.
Ask a clarifying question that will help you determine if any member of the user’s household is eligible for benefits as efficiently as possible. Only ask about one fact at a time.
C.4 “Predict Benefits Eligibility” for CoT Prompt
Eligibility requirements: {eligibility_requirements}.
Is the information sufficient to determine whether any member of the user’s household is eligible for all programs? Think through your reasoning out loud. Then answer with True or False.
C.5 “Predict Benefits Reasoning” for CoT Prompt
Eligibility: {eligibility_requirements}.
Predict the programs for which any member of the user’s household is eligible. Return only a boolean array of length {num_programs}, e.g. {example_array}, where the value at index ‘i‘ is true iff the user is eligible for program ‘i‘. Only return the array. Do not return anything else in the response. If a user’s eligibility is unclear, make your best guess. Think through your reasoning out loud.
C.6 “Predict Benefits Constrained” for CoT Prompt
Reasoning: {reasoning}.
Using the reasoning above, predict the programs for which any member of the user’s household is eligible. Output a boolean array of length {num_programs}, e.g. {example_array}, where the value at index ‘i‘ is true iff the user is eligible for program ‘i‘. If a user’s eligibility is unclear, make your best guess.
C.7 “Predict Clarifying Questions” for ReAct Chain-of-Thought Prompt
Eligibility: {eligibility_requirements}.
Ask a clarifying question that will help you determine if any member of the user’s household is eligible for benefits as efficiently as possible. Only ask about one fact at a time. Think through your reasoning out loud, then state your question after a colon, e.g. Question: What is the user’s age?
C.8 “Generate Checker” Prompt
{attempt_no}
Eligibility Requirements: {eligibility_requirement}
Write a python function called check_eligibility that takes a dictionary hh containing relevant information and determines user eligibility. hh is a special dictionary connected to a language model that is conversing with the user. Any time it does not contain a key, it will determine that information from the user. As a result here are some requirements for interacting with hh:
  • DO NOT use dict.get() anywhere in the code. Key errors will be handled elsewhere.
  • Do not use default values.
  • Do not use any f-strings, curly brackets, or dynamically generated strings in your keys.
  • Use only literal strings in keys.
  • Do not use try-except blocks.
  • If you need to access data for individuals (rather than the household as a whole) you can use integer indexing. hh[0] is the head of the household.
check_eligibility returns a bool. All keys and values of hh are strings. If you write helper functions, keep them inside the check_eligibility function. Make your code as detailed as possible capturing every edge case. Remember that the household may have no relevant members, so be sure to ask about the composition of the household. For example, for childcare programs, check that the household has at least one child. After each new lookup in hh, write a comment suggesting a question to ask.
The following is a set of preexisting keys and values in the hh dictionary; take care not to duplicate them.
{preexisting_keys}
Avoid using int() and use float() instead. Do not provide anything besides code in your response. Do not use input for user input.
C.9 “Get Type” Prompt
Context: {eligibility_requirements}
Code: {code}
Target key: {key}
Question: Given the code and context above, what do you expect {key} to be an integer, a float, or one choice from a set of strings? Return ONLY int, float, or choice.
C.10 “Get Values” Prompt
Context: {eligibility_requirements}
Code: {code}
Target key: {key}
Question: Given the code and context above, what are the possible values of {key}? Return ONLY the list of possible values in a list of strings. For example, return ["a", "b", "c"].
C.11 “Extract Values from Answer” Prompt
Context: {eligibility_requirements}
Line: ‘‘‘{line}‘‘‘
We need to extract the value of {key} from the following dialog:
Question: {cq} Answer: {answer}
What should we set as the value of {key}? Return ONLY the value.
C.12 “Key Error” Prompt
Context: {eligibility_requirements}
Line: ‘‘‘{line}‘‘‘
We need to determine what value of {key} should be stored in the hh dictionary. Ask a question to the user that would get this value. For example, for age_i, ask “What is the age of person i?”. Return ONLY the question.
Benefits Program Positive Count Negative Count Percentage True (%)
FamilyHomelessnessAndEvictionPreventionSupplement 4 3 57.14
WorkforceoneCareerCenters 1 1 50.00
SilverCorps 1 1 50.00
AdultProtectiveServices 1 2 33.33
DisabilityRentIncreaseExemption 8 7 53.33
ChildTaxCredit 1 4 20.00
SeniorCitizenHomeownersExemption 1 9 10.00
InfantToddlerPrograms 3 6 33.33
LearnEarn 7 1 87.50
DisabledHomeownersExemption 1 6 14.29
PreKForAll 1 1 50.00
JobsPlus 1 1 50.00
HeadStart 6 2 75.00
KindergartenAndElementarySchool 1 1 50.00
CoolingAssistanceBenefit 4 5 44.44
HomeEnergyAssistanceProgram 4 3 57.14
VeteransAffairsSupportedHousing 1 1 50.00
NYCFreeTaxPrep 2 1 66.67
FamilyPlanningBenefitProgram 1 6 14.29
ChildrenAndYouthWithSpecialHealthCareNeeds 3 2 60.00
EnhancedSchoolTaxReliefProgram 0 2 0.00
SummerMeals 1 1 50.00
TrainEarn 6 3 66.67
NYCFinancialEmpowermentCenters 1 1 50.00
NYCHAPublicHousing 3 5 37.50
ChildAndDependentCareTaxCredit 4 4 50.00
ChildCareVouchers 8 2 80.00
HIVAIDSServicesAdministration 1 1 50.00
BigAppleConnect 1 1 50.00
OfficeOfChildSupportServices 1 3 25.00
BeaconPrograms 1 1 50.00
SafeAndSickLeave 1 5 16.67
NYSUnemploymentInsurance 1 1 50.00
FamilyTypeHomesForAdults 1 6 14.29
EarnedIncomeTaxCredit 1 6 14.29
Homebase 1 1 50.00
HomeFirstDownPaymentAssistance 2 1 66.67
HighSchool 1 1 50.00
SeniorCitizenRentIncreaseExemption 1 1 50.00
AccessARideParatransitService 2 1 66.67
TextTwoWork 4 1 80.00
TheEarlyInterventionProgram 1 1 50.00
EarlyHeadStart 2 4 33.33
Lifeline 6 1 85.71
IDNYC 1 1 50.00
NYSPaidFamilyLeave 2 1 66.67
COVIDnineteenFuneralAssistance 1 1 50.00
SchoolAgeAndEarlyChildhoodFamilyAndCommunityEngagementFACECenters 1 1 50.00
FairFaresNYC 1 2 33.33
NYCYouthHealth 1 1 50.00
NewbornHomeVisitingProgram 4 2 66.67
AcceleratedStudyInAssociatePrograms 1 1 50.00
STEMMattersNYC 1 1 50.00
CommoditySupplementalFoodProgram 1 2 33.33
CareerAndTechnicalEducation 1 1 50.00
NYCHAResidentEconomicEmpowermentAndSustainability 1 1 50.00
OutpatientTreatmentServices 1 1 50.00
CUNYFatherhoodAcademy 1 3 25.00
SummerYouthEmploymentProgram 1 1 50.00
ThreeK 1 1 50.00
MedicaidForPregnantWomen 1 2 33.33
ActionNYC 1 1 50.00
FamilyResourceCenters 2 1 66.67
NYCCare 1 1 50.00
PrimaryAndPreventiveHealthCare 1 1 50.00
NYCTenantResourcePortal 1 1 50.00
OlderAdultEmploymentProgram 1 1 50.00
NYCLaddersForLeaders 1 1 50.00
CornerstonePrograms 1 1 50.00
ComprehensiveAfterSchoolSystemOfNYC 1 1 50.00
WeSpeakNYC 1 1 50.00
NYCMitchellLama 1 0 100.00
CUNYStart 1 1 50.00
NYCNurseFamilyPartnership 1 1 50.00
MiddleSchool 1 1 50.00
AdvanceEarn 1 1 50.00
SectionEightHousingChoiceVoucherProgram 1 0 100.00
NYCYouthLeadershipCouncils 1 1 50.00
ChildHealthPlusAndChildrensMedicaid 1 2 33.33
VeteransPropertyTaxExemption 1 1 50.00
FamilyAssessmentProgram 1 1 50.00
BasicSchoolTaxReliefProgram 1 0 100.00
Total 146 159 47.88
Table 3: Benefits Program-wise Positive/Negative Counts and Percentages for the Diversity Dataset
Benefits Program Positive Count Negative Count Percentage True (%)
AdultProtectiveServices 0 3 0.00
HomeEnergyAssistanceProgram 0 3 0.00
MiddleSchool 0 3 0.00
NYCFreeTaxPrep 0 3 0.00
NYSPaidFamilyLeave 1 2 33.33
NYSUnemploymentInsurance 0 3 0.00
WorkforceoneCareerCenters 1 2 33.33
FamilyTypeHomesForAdults 2 1 66.67
HeadStart 0 3 0.00
NYCFinancialEmpowermentCenters 3 0 100.00
TextTwoWork 3 0 100.00
ThreeK 0 3 0.00
CornerstonePrograms 0 3 0.00
JobsPlus 0 3 0.00
CommoditySupplementalFoodProgram 0 3 0.00
NYCCare 0 3 0.00
SilverCorps 0 3 0.00
SummerMeals 1 2 33.33
PrimaryAndPreventiveHealthCare 1 2 33.33
IDNYC 3 0 100.00
NYCYouthLeadershipCouncils 0 3 0.00
Homebase 1 2 33.33
NYCMitchellLama 2 1 66.67
NYCNurseFamilyPartnership 0 3 0.00
AdvanceEarn 0 3 0.00
BeaconPrograms 1 2 33.33
ChildHealthPlusAndChildrensMedicaid 0 3 0.00
BasicSchoolTaxReliefProgram 0 3 0.00
TheEarlyInterventionProgram 0 3 0.00
AccessARideParatransitService 3 0 100.00
ChildAndDependentCareTaxCredit 0 3 0.00
FamilyResourceCenters 0 3 0.00
InfantToddlerPrograms 0 3 0.00
NYCTenantResourcePortal 3 0 100.00
NYCYouthHealth 1 2 33.33
DisabledHomeownersExemption 0 3 0.00
OutpatientTreatmentServices 0 3 0.00
STEMMattersNYC 0 3 0.00
SeniorCitizenHomeownersExemption 0 3 0.00
CareerAndTechnicalEducation 0 3 0.00
NewbornHomeVisitingProgram 0 3 0.00
SchoolAgeAndEarlyChildhoodFamilyAndCommunityEngagementFACECenters 0 3 0.00
BigAppleConnect 0 3 0.00
CUNYFatherhoodAcademy 0 3 0.00
HomeFirstDownPaymentAssistance 0 3 0.00
DisabilityRentIncreaseExemption 0 3 0.00
KindergartenAndElementarySchool 0 3 0.00
EarnedIncomeTaxCredit 1 2 33.33
HIVAIDSServicesAdministration 0 3 0.00
OlderAdultEmploymentProgram 0 3 0.00
FamilyHomelessnessAndEvictionPreventionSupplement 0 3 0.00
ChildCareVouchers 1 2 33.33
ComprehensiveAfterSchoolSystemOfNYC 1 2 33.33
COVIDnineteenFuneralAssistance 0 3 0.00
TrainEarn 0 3 0.00
LearnEarn 0 3 0.00
SectionEightHousingChoiceVoucherProgram 0 3 0.00
CoolingAssistanceBenefit 0 3 0.00
MedicaidForPregnantWomen 0 3 0.00
SummerYouthEmploymentProgram 1 2 33.33
FairFaresNYC 0 3 0.00
PreKForAll 0 3 0.00
ChildrenAndYouthWithSpecialHealthCareNeeds 1 2 33.33
CUNYStart 2 1 66.67
NYCLaddersForLeaders 0 3 0.00
FamilyAssessmentProgram 1 2 33.33
FamilyPlanningBenefitProgram 0 3 0.00
NYCHAPublicHousing 0 3 0.00
SafeAndSickLeave 2 1 66.67
WeSpeakNYC 1 2 33.33
VeteransAffairsSupportedHousing 0 3 0.00
NYCHAResidentEconomicEmpowermentAndSustainability 0 3 0.00
SeniorCitizenRentIncreaseExemption 0 3 0.00
AcceleratedStudyInAssociatePrograms 0 3 0.00
EnhancedSchoolTaxReliefProgram 0 3 0.00
EarlyHeadStart 0 3 0.00
ActionNYC 0 3 0.00
Lifeline 1 2 33.33
VeteransPropertyTaxExemption 0 3 0.00
HighSchool 0 3 0.00
OfficeOfChildSupportServices 0 3 0.00
ChildTaxCredit 0 3 0.00
Total 38 208 15.42
Table 4: Benefits Program-wise Positive/Negative Counts and Percentages for the Representative Dataset
Appendix D Instructions to User Survey Participants
Please select five programs in which you are interested. The chatbot will help you determine your eligibility for them. Please answer the questions with respect to your own household. You may end the study at any time. Your responses will not be stored. Only the final model prediction and your feedback will be used.
Participants were students and a New York City university recruited in person and compensated up to \$5 USD.
Appendix E Constrained Generation Implementation Details
For constrained generation, when available (e.g., for OpenAI models) we use tool use JSON formatting to ensure constraints are satisfied. Where the constrains cannot be clearly specified in the API, we check whether the output conforms to a validation function (e.g., with regular expressions) and regenerate it if it does not. For hugging face models, we use the Outlines library choices and regex functions in most cases Willard and Louf (2023). For ReAct-style prompting, we first instruct the model to think through their reasoning. Then, in a second prompt, we prompt the model to output its answer in the specified form based on the thoughts.